# Game-Theoretic Coordination For Time-Critical Missions of UAV Systems

### v1

@@ III Main Approach
In this section, we prove the existence and exponential stability of the solution to Problem 1. Furthermore, we introduce a MPC based algorithm that approximates the solution of Problem 1.
@@ III-A Existence of Solution and Exponential Stability
Under ideal system conditions, we isolate and analyze the core features of the game-theoretic approach. More precisely, to establish the existence and exponential stability of the solution to Problem 1, we first prove the existence and exponential stability of the solution to the corresponding unconstrained problem. Then, we determine the parameter $\alpha$ (discount rate) such that the solution to the unconstrained problem is also a solution to the constrained problem. We then conclude that, for the chosen $\alpha$, Problem 1 has a solution and that the solution is exponentially stable.
@@
$\begin{split}\int_{0}^{\infty}e^{-\alpha t}\Bigg{(}\dot{\gamma}_{i}^{2}+\sum_{j=1}^{N}(\gamma_{i}-\gamma_{j})^{2}+\ddot{\gamma}_{i}^{2}\Bigg{)}{\rm d}t\end{split}$ (14)
over $\gamma_{i}\in\mathcal{B}^{0}_{i}=\{\gamma_{i}\in H^{1}_{w,\alpha}((0,\infty)):(\gamma_{i}+t)\in\mathcal{A}^{0}_{i}\}$.
The following result proves the existence and exponential stability of the solution to Problem 2 and provides an explicit expression.
@@
Let $\alpha>0$. Then, there exists $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0}_{j}$ solving Problem 2. Moreover, the solution has the following explicit form
$\gamma_{i}^{*}(t)=H^{1}_{i}+H^{3}_{i}e^{\mu_{3}t}+e^{\mu_{1}t}\Big{(}C_{i}^{1}\cos(\nu_{1}t)+C_{i}^{2}\sin(\nu_{1}t)\Big{)},$ (15)
where $\mu_{1},\mu_{3}<0$, and the constants $H^{1}_{i},H_{i}^{3},C_{i}^{1},\nu_{1},C_{i}^{2}$ only depend on $\alpha$ and initial conditions.
@@
Here, we present the general idea of the proof; for more details see Appendix IX.
Suppose that $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0}_{j}$ is a solution to Problem 2. Then, by the definition of Nash equilibrium and the convexity of the integrand of (14) it follows that $\gamma_{i}^{*}$ is the unique minimizer of the following optimization problem
[…]
@@
$\gamma_{i}^{(4)*}-2\alpha\dddot{\gamma}_{i}^{*}+(\alpha^{2}-1)\ddot{\gamma}_{i}^{*}+\alpha\dot{\gamma}_{i}^{*}+\sum_{j=1}^{N}(\gamma_{i}^{*}-\gamma_{j}^{*})=0,$ (17)
@@
$\begin{split}&\lim_{T\to\infty}e^{-\alpha T}\ddot{\gamma}_{i}^{*}(T)=0,\\ &\lim_{T\to\infty}e^{-\alpha T}\Big{(}\dot{\gamma_{i}}^{*}(T)+\alpha\ddot{\gamma}_{i}^{*}(T)-\dddot{\gamma}_{i}^{*}(T)\Big{)}=0.\end{split}$ (18)
Using (18) and the boundary conditions (10), we solve the system of Euler-Lagrange equations in (17) and obtain (15).
To prove the existence of the solution to Problem 2, we repeat the steps above in the opposite direction. Specifically, since $\gamma_{i}^{*}$ in (15) is solving the Euler-Lagrange equation in (17), and the variational problem in (16) is quadratic, we have that $\gamma_{i}^{*}$ is the minimizer of (16). Subsequently, $\bm{\gamma}^{*}$ is a Nash equilibrium of Problem 2. ∎
Next, we prove the existence and exponential stability of the solution to the constrained problem, i.e. Problem 1.
@@
For any initial condition, $\gamma^{0}=(\gamma_{1}^{0},\dots,\gamma_{N}^{0})$, and physical constraints on the UAVs ($V^{i}_{1}$ and $V_{2}^{i}$), there exists $\alpha>0$ such that Problem 1 has a solution, and it is exponentially stable.
@@
In the detailed proof of Proposition 2 (see Appendix IX), we keep track of the exact dependence of the constants $H^{1}_{i},H_{i}^{3},\mu_{1},\mu_{3},\nu_{1},C_{i}^{1},C^{2}_{i}$ on the parameter $\alpha$. Using these dependencies, we have
$\begin{split}H^{1}_{i}=\gamma_{i}^{0}+O(\tfrac{1}{\alpha^{2}}),\quad H_{i}^{3}=O(\tfrac{1}{\alpha^{3}}),\quad C^{1}_{i}=O(\tfrac{1}{\alpha^{2}}),\\ C^{2}_{i}=O(\tfrac{1}{\alpha}),\quad\mu_{3}=O(\tfrac{1}{\alpha}),\quad\mu_{1}=O(\tfrac{1}{\alpha}),\quad\nu_{1}=O(\tfrac{1}{\alpha}),\end{split}$ (19)
as $\alpha\to\infty$. On the other hand, the explicit solution to Problem 2 (see (15)) is a combination of uniformly bounded functions in $t$
$\cos(\nu_{1}t),\quad\sin(\nu_{1}t),\quad e^{\mu_{1}t},\quad e^{\mu_{3}t}.$
This with (19) implies that taking large enough $\alpha$ for the explicit solution to Problem 2, we obtain $(\gamma_{i}^{*}-t)\in\mathcal{A}^{2,\alpha}_{i}$. Therefore, $(\bm{\gamma}^{*}-t)$ is a solution to the constrained problem, Problem 1, as well. By Proposition 2, it follows that the solution is exponentially stable. ∎
@@ III-B Algorithm
In this part, we develop an algorithm using Problem 1 that leads to time coordination among UAVs. The algorithm uses MPC, resulting in a finite-dimensional optimization problem, which serves as an approximation to the solution to Problem 1, while accommodating real-world challenges such as communication constraints, path-following errors and dynamic environments.
In the formulation of Problem 1, the communication network is static and completely connected; i.e. each UAV is connected to all others involved in the mission. The proposed algorithm relaxes these assumptions. We assume that the communication network remains static during each MPC step but may change between steps. The communication network change is incorporated into the cost function. Furthermore, the algorithm considers the path-following errors in the constraints that may arise when each UAV follows its desired trajectory operating in an uncertain and dynamic environment.
@@
for $k=1,2,\dots$. The optimization variables are
$y^{k}_{i}=[s^{k}_{i},\ell^{k}_{i},u^{k}_{i}],$
where
$\begin{split}s^{k}_{i}=[s^{k}_{i0},s^{k}_{i1},\dots,s^{k}_{iK}],\\ \ell^{k}_{i}=[\ell^{k}_{i0},\ell^{k}_{i1},\dots,\ell^{k}_{iK}],\\ u^{k}_{i}=[u^{k}_{i0},u^{k}_{i1},\dots,u^{k}_{iK-1}].\end{split}$
@@
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i}\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}+F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i}\right)+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2},$ (20)
with
$\bar{s}^{k}_{-i}=\left[\bar{s}^{k}_{1},\dots,\bar{s}^{k}_{i-1},\bar{s}^{k}_{i+1},\dots,\bar{s}^{k}_{N}\right]$ (21)
and
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i}\right)=\sum_{j\in\mathcal{N}_{ik}}\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$ (22)
where $\mathcal{N}_{ik}$ denotes the neighborhood of $i^{\rm th}$ UAV, which is the set of UAVs that communicate with $i^{\rm th}$ one at time $t_{k}$.
All terms of the cost function in (20) have been discussed in the continuous case; see Section II-C. The newly added correction term $\alpha_{i}^{k}$ appears in initial conditions and takes into account the path-following error:
$\alpha_{i}^{k}=\alpha_{i}^{k}(x_{i}(t_{k}))=\beta\dfrac{e_{x,i}(t_{k})^{T}\dot{x}_{\gamma,i}(t_{k})}{\left|\left|\dot{x}_{\gamma,i}(t_{k})\right|\right|+\delta},$ (23)
where $\beta$ and $\delta$ are positive parameters. The correction term $\alpha_{i}^{k}$ is positive if the UAV’s actual location projection on the desired trajectory is ahead of the desired position $x_{\gamma,i}(t_{k})$ and negative otherwise. Including $\alpha_{i}^{k}$ in initial conditions causes the UAV to slow down to reduce the forward overshoot and to accelerate to catch up with the desired trajectory.
At next time instance $t_{k+1}$ each UAV shares the computed $[s^{k}_{i1},\dots,s^{k}_{iK}]$ (solution to Problem 3) with the UAVs that are from its neighborhood $\mathcal{N}_{ik+1}$ at time instance $t_{k+1}$. For the $j^{\rm th}$ UAV from $\mathcal{N}_{ik+1}$, we have $\bar{s}^{k+1}_{j\tau}=s^{k}_{j\tau+1}$, $\tau=1,\dots,K-1$. The algorithm starts at time $t_{1}$ ($k=1$) with $s_{i0}^{1}=\gamma_{i}^{0}$, $l_{i0}^{1}=\dot{\gamma}_{i}^{0}$ and $\bar{s}_{j\tau}^{1}=\gamma_{j}^{0}+\tau h$, $\tau=1,\dots,K$, which are known due to initial information sharing between neighboring UAVs. Eventually, as approximations to $\gamma_{i}(t)$ and to its first-order and second-order derivatives, we set
@@
The second term of the cost function (20) shows that the change in the communication can significantly impact the virtual time, particularly its second-order derivative. To mitigate and smoothen these effects, mainly in scenarios where communication depends on the distance between UAVs, we introduce the smoothing function:
$\phi\left(z,p_{1},p_{2}\right)=\begin{cases}1,&z<p_{1},\\ \eta(z),&p_{1}\leqslant z\leqslant p_{2},\\ 0,&p_{2}<z,\end{cases}$ (24)
where $p_{1}<p_{2}$ are positive parameters and $0\leqslant\eta(z)\leqslant 1$, such that $\phi\in C^{2}({\mathbb{R}}^{+})$ for any $p_{1}<p_{2}$.
The function $\phi$ quantifies the quality of communication between UAVs based on their relative distances. Communication is excellent if the distance between two UAVs is less than $p_{1}$. In contrast, communication is lost if the distance exceeds $p_{2}$. For distances in the intermediate range $\left[p_{1},p_{2}\right]$, the quality of communication smoothly transitions between these two extremes. The modified cost function is
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}\\ +F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2}$ (25)
with
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=\sum_{j\neq i}&\phi\left(d_{i,j}^{k,0},a,b\right)\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$
where
$d_{i,j}^{k,\tau}=\left|\left|x_{d,i}\left(s^{k}_{i\tau}\right)-x_{d,j}\left(\bar{s}^{k}_{j\tau}\right)\right|\right|.$
Although the smoothing function is not convex, the modified cost function is quadratic because the function $\phi$ is treated as a constant in the cost function for each MPC step. Since it also does not impact the constraints, Problem 3 remains a quadratic optimization problem.
Remark 2. The formulation of Problem 3 is agile and can be modified to achieve additional goals; for example, the cost function can be augmented with additional terms that ensure collision avoidance:
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2}\\ +F_{i}^{ca}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)+F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right).$ (26)

### v2

@@ III Main Approach
In this section, we prove the existence and exponential stability of the solution to Problem 1. Furthermore, we introduce an MPC based algorithm that approximates the solution of Problem 1.
@@ III-A Existence of Solution and Exponential Stability
Under ideal system conditions, we isolate and analyze the core features of the game-theoretic approach. More precisely, to establish the existence and exponential stability of the solution to Problem 1, we first prove the existence and exponential stability of the solution to the corresponding unconstrained problem. Then, we determine the parameter $\alpha$ (discount rate) such that the solution to the unconstrained problem is also a solution to the constrained problem. We then conclude that, for the chosen $\alpha$, Problem 1 has a solution that exhibits exponential convergence with respect to the equilibrium trajectory.
@@
$I^{\alpha}_{\bm{\gamma}}[\gamma_{i}]=\int_{0}^{\infty}e^{-\alpha t}\Big(w_{1}\dot{\gamma}_{i}^{2}+\tfrac{w_{2}}{N}\sum_{j=1}^{N}(\gamma_{i}-\gamma_{j})^{2}+w_{3}\ddot{\gamma}_{i}^{2}\Big){\rm d}t$ (12)
over $\gamma_{i}\in\mathcal{B}^{0}_{i}:=\{\gamma_{i}\in H^{2}_{w,\alpha}((0,\infty)):(\gamma_{i}+t)\in\mathcal{A}^{0}_{i}\}$.
Note that in $\gamma_{i}(t)$ the argument $t$ refers to the physical time of the UAV. This is distinct from the variable $t$ used inside the optimization problem (12), which serves as a dummy variable of integration in the optimization. The following result proves the existence and exponential stability of the solution to Problem 2 and provides an explicit expression.
@@
Let $\alpha,w_{3}>0$, $w_{1},w_{2}\geqslant 0$, and $w_{1}+w_{2}+w_{3}=1$ and suppose that Assumptions 1-2 hold. Then, there exists a unique $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0,\alpha}_{j}$ solving Problem 2. Moreover, letting $W:=w_{1}^{2}-4w_{2}w_{3}$ and $z_{i}^{r}(t):=H_{1i}^{r}+H_{3i}^{r}e^{\mu_{3}t}$ (for $r\in\{0,+,-\}$), the solution takes the following explicit form:
$\begin{split}\gamma_{i}^{*}(t)=z_{i}^{0}(t)+C^{0}_{1i}e^{\mu_{1}^{0}t}+C^{0}_{2i}te^{\mu^{0}t},\end{split}$ (13)
when $W=0$;
$\begin{split}\gamma_{i}^{*}(t)=z_{i}^{+}(t)+C^{+}_{1i}e^{\mu^{+}_{1}t}+C^{+}_{2i}e^{\mu^{+}_{2}t},\end{split}$ (14)
when $W>0$; and
$\gamma_{i}^{*}(t)=z_{i}^{-}(t)+e^{\mu^{-}_{1}t}(C^{-}_{i1}\cos(\nu_{1}t)+C^{-}_{i2}\sin(\nu_{1}t)),$ (15)
when $W<0$. The constants $\mu_{3}$, $\mu_{1}^{r}$, $\mu_{2}^{+}$, $H_{1i}^{r}$, $H_{2i}^{r}$, $C_{i1}^{r}$, $C_{i2}^{r}$, and $\nu_{1}$ (for $r\in\{0,+,-\}$) depend only on $\alpha$, $w_{1}$, $w_{2}$, $w_{3}$, and the initial conditions, and satisfy $\mu_{3},\mu_{1}^{r},\mu_{2}^{+}<0$.
@@
Here, we present the general idea of the proof; for more details see Appendix VI. Suppose that $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0}_{j}$ is a solution to Problem 2. Then, by the definition of Nash equilibrium and the convexity of the integrand of (12) it follows that $\gamma_{i}^{*}$ is the unique minimizer of the following optimization problem
$\begin{split}I^{\alpha}_{\bm{\gamma}^{*}}[\gamma_{i}^{*}]&=\min_{\gamma_{i}\in\mathcal{B}^{0}_{i}}I^{\alpha}_{\bm{\gamma}^{*}}[\gamma_{i}].\end{split}$ (16)
@@
$\gamma_{i}^{(4)*}-2\alpha\dddot{\gamma}_{i}^{*}+(\alpha^{2}-\tfrac{w_{1}}{w_{3}})\ddot{\gamma}_{i}^{*}+\tfrac{\alpha w_{1}}{w_{3}}\dot{\gamma}_{i}^{*}+\tfrac{w_{2}}{w_{3}N}\sum_{j=1}^{N}(\gamma_{i}^{*}-\gamma_{j}^{*})=0,$ (17)
@@
$\begin{split}&\lim_{T\to\infty}e^{-\alpha T}\ddot{\gamma}_{i}^{*}(T)=0,\\ &\lim_{T\to\infty}e^{-\alpha T}(\dot{\gamma_{i}}^{*}(T)+\alpha\ddot{\gamma}_{i}^{*}(T)-\dddot{\gamma}_{i}^{*}(T))=0.\end{split}$ (18)
Using (18) together with the boundary conditions in $\mathcal{B}_{i}^{0}$, we solve the system of Euler–Lagrange equations in (17) and find the unique solution. Depending on the sign of the parameter $W$, we obtain (13), (14), or (15), respectively.
To prove the existence of the solution to Problem 2, we note that since $\gamma_{i}^{*}$ solves the Euler-Lagrange equation in (17), and the variational problem in (16) is quadratic, we have that $\gamma_{i}^{*}$ is the minimizer of (16). Subsequently, $\bm{\gamma}^{*}$ is a Nash equilibrium of Problem 2. ∎
Next, we prove the existence and exponential convergence of the solution to the constrained problem, i.e. Problem 1.
@@
Let the weights satisfy one of the following conditions: $w_{1}^{2}-4w_{2}w_{3}<0$ or $w_{1}^{2}-4w_{2}w_{3}\geqslant 0$ with $w_{1}=w_{3}O(\alpha)$. Suppose that Assumptions 1-2 hold. Then, for any initial condition $\gamma^{0}=(\gamma_{1}^{0},\dots,\gamma_{N}^{0})$ and given the physical constraints of the UAVs (denoted by $V_{1}^{i}$ and $V_{2}^{i}$), there exists $\alpha>0$ such that Problem 1 admits a unique solution, and the corresponding trajectory exhibits exponential convergence to the equilibrium trajectory.
@@
In the detailed proof of Proposition 3 (see Appendix VI), we explicitly track the dependence of the constants $\mu_{3}$, $\mu_{1}^{r}$, $\mu_{2}^{+}$, $H_{1i}^{r}$, $H_{2i}^{r}$, $C_{i1}^{r}$, $C_{i2}^{r}$, and $\nu_{1}$ (for $r\in\{0,+,-\}$) on the parameters $\alpha$, $w_{1}$, $w_{2}$, and $w_{3}$. Using these dependencies as $\alpha\to\infty$, we have
$H^{r}_{1i}=\gamma_{i}^{0}+O(\tfrac{1}{\alpha}),\mu_{3}=O(\tfrac{1}{\alpha}),\mu^{r}_{1}=O(\tfrac{1}{\alpha}),\mu^{+}_{2}=O(\tfrac{1}{\alpha})$
$\nu_{1}=O(\tfrac{1}{\alpha}),H^{r}_{i3}=O(\tfrac{1}{\alpha}),C^{r}_{1i}=O(\tfrac{1}{\alpha}),C^{r}_{2i}=O(\tfrac{1}{\alpha}),$ (19)
where $r\in\{0,+,-\}$. On the other hand, the explicit solution to Problem 2 (see (13), (14), (15)) is a combination of uniformly bounded functions in $t$: $\cos(\nu_{1}t),\quad\sin(\nu_{1}t),\quad e^{\mu_{1}t},\quad e^{\mu_{3}t}$. This with (19) implies that taking large enough $\alpha$ for the explicit solution to Problem 2, we obtain $(\gamma_{i}^{*}+t)\in\mathcal{A}^{2,\alpha}_{i}$. Therefore, $(\bm{\gamma}^{*}+t)$ is a solution to the constrained problem, Problem 1, as well. By Proposition 2, it follows that the solution exhibits exponential convergence with respect to the equilibrium trajectory. ∎
@@ III-B Algorithm
In this part, we develop an algorithm based on Problem 1 to achieve time coordination among UAVs. The algorithm employs the idea of MPC, as a natural discrete-time continuation of the original optimization problem 1 (see Fig. 2). Beyond this conceptual consistency, the MPC formulation offers several practical advantages under non-ideal conditions, such as variations in network topology and path-following errors. In particular, MPC provides a systematic mechanism for enforcing input and state constraints at each time step while optimizing over a finite prediction horizon, which is an essential capability for UAV coordination, where actuation limits, such as constraints on velocity and acceleration, must be satisfied. Moreover, MPC formulation enables the accommodation of time-varying objectives of individual UAVs and autonomous adaptation to environmental changes and external disturbances. For instance, the effectiveness of this feature is illustrated through the demonstrated collision avoidance behaviour. To clarify the relationship between the continuous-time coordination problem (Problem 1) and the discrete-time MPC formulation presented in this section (Problem 3), we present a flow chart in Fig. 2.
Problem 1:Continuous-time coordinationDiscretization:Infinite-horizon formulationProblem 3:MPC optimizationDiscretize time, reformulate cost/constraintsReceding horizon, real-time solution
Figure 2: Flow chart showing the transition from the continuous-time coordination problem (Problem 1) to the discrete-time MPC problem (Problem 3).
In the formulation of Problem 1, the communication network is static and completely connected; i.e., each UAV is connected to all others involved in the mission. The proposed algorithm relaxes these assumptions. We assume that the communication network remains static during each MPC step but may change between steps. The communication network change is incorporated into the cost function. Furthermore, the algorithm considers the path-following errors in the constraints that may arise when each UAV follows its desired trajectory operating in an uncertain and dynamic environment.
@@
for $k=1,2,\dots$. The optimization variables are $y^{k}_{i}=[s^{k}_{i},\ell^{k}_{i},u^{k}_{i}]$, where $s^{k}_{i}=[s^{k}_{i0},s^{k}_{i1},\dots,s^{k}_{iK}]$ and
$\begin{split}\ell^{k}_{i}=[\ell^{k}_{i0},\ell^{k}_{i1},\dots,\ell^{k}_{iK}],\quad u^{k}_{i}=[u^{k}_{i0},u^{k}_{i1},\dots,u^{k}_{iK-1}].\end{split}$
@@
$J_{i}(y^{k}_{i},\bar{s}^{k}_{-i})=\sum_{\tau=1}^{K}(\ell^{k}_{i\tau}-1)^{2}+F_{i}(s^{k}_{i},\bar{s}^{k}_{-i})+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2},$ (20)
with $\bar{s}^{k}_{-i}=\left[\bar{s}^{k}_{1},\dots,\bar{s}^{k}_{i-1},\bar{s}^{k}_{i+1},\dots,\bar{s}^{k}_{N}\right]$ and $F_{i}(s^{k}_{i},\bar{s}^{k}_{-i})=\sum_{j\in\mathcal{N}_{ik}}\sum_{\tau=1}^{K}(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau})^{2}$, where $\mathcal{N}_{ik}$ denotes the neighborhood of $i^{\rm th}$ UAV, which is the set of UAVs that communicate with $i^{\rm th}$ one at a time $t_{k}$. All terms of the cost function in (20) have been discussed in the continuous case; see Section II-C. The newly added correction term $\alpha_{i}^{k}$ appears in initial conditions and takes into account the path-following error:
[…]
where $\beta$ and $\delta$ are positive parameters and $x_{i}(t_{k})$ is the $i^{\rm th}$ UAV’s actual position. The correction term $\alpha_{i}^{k}$ is negative if the UAV’s actual location projection on the desired trajectory is ahead of the desired position $x_{\gamma,i}(t_{k})$ and positive otherwise. Including $\alpha_{i}^{k}$ in initial conditions causes the UAV to slow down to reduce the forward overshoot and to accelerate to catch up with the desired trajectory.
Remark 5
Unlike the ideal setting of Problem 1 (see Assumption 2 and Remark 1), Problem 3 explicitly accounts for path-following errors. Such errors may arise when the trajectory reparameterized by the virtual time—subject only to the linear constraints in (6) and (7)—is not dynamically feasible. These effects are incorporated through the correction term $\alpha_{i}$. Provided that each UAV employs a sufficiently accurate low-level path-following controller, the proposed coordination method remains effective. This is validated by the simulation results in Section IV-A3.
The parameter $\delta$ in the correction term $\alpha_{i}^{k}$ is introduced to prevent division by zero; it can be set as a positive constant close to $1$ without affecting the overall performance. The parameter $\beta$ regulates the influence of the path-following error on the coordination process. Increasing $\beta$ accelerates time coordination but may cause constraint violations if chosen excessively large.
At next time instance $t_{k+1}$ each UAV shares the computed $s^{k}_{i1},\dots,s^{k}_{iK}]$ (solution to Problem 3) with the UAVs that are from its neighborhood $\mathcal{N}_{ik+1}$ at time instance $t_{k+1}$. For the $j^{\rm th}$ UAV from $\mathcal{N}_{ik+1}$, we have $\bar{s}^{k+1}_{j\tau}=s^{k}_{j\tau+1}$, $\tau=1,\dots,K-1$. The algorithm starts at time $t_{1}$ ($k=1$) with $s_{i0}^{1}=\gamma_{i}^{0}$, $l_{i0}^{1}=\dot{\gamma}_{i}^{0}$ and $\bar{s}_{j\tau}^{1}=\gamma_{j}^{0}+\tau h$, $\tau=1,\dots,K$, which are known due to initial information sharing between neighboring UAVs. Eventually, as approximations to $\gamma_{i}(t)$ and to its first-order and second-order derivatives, we set
@@
The second term of the cost function (20) shows that the change in the communication can significantly impact the virtual time, particularly its second-order derivative. To mitigate and smoothen these effects, mainly in scenarios where communication depends on the distance between UAVs, we introduce the smoothing function: $\phi\left(z,p_{1},p_{2}\right)=1$, when $z<p_{1}$, $\phi\left(z,p_{1},p_{2}\right)=\eta(z)$, when $p_{1}\leqslant z\leqslant p_{2}$ and $\phi\left(z,p_{1},p_{2}\right)=0$, when $p_{2}<z$. Here, $p_{1}<p_{2}$ are positive parameters and $0\leqslant\eta(z)\leqslant 1$, such that $\phi\in C^{2}({\mathbb{R}}^{+})$ for any $p_{1}<p_{2}$. The function $\phi$ quantifies the quality of communication between UAVs based on their relative distances. Communication is excellent if the distance between two UAVs is less than $p_{1}$. In contrast, communication is lost if the distance exceeds $p_{2}$. For distances in the intermediate range $\left[p_{1},p_{2}\right]$, the quality of communication smoothly transitions between these two extremes. The modified cost function is
$J_{i}(y^{k}_{i},\bar{s}^{k}_{-i})=\sum_{\tau=1}^{K}(\ell^{k}_{i\tau}-1)^{2}+F_{i}^{\phi}(s^{k}_{i},\bar{s}^{k}_{-i})+\sum_{\tau=0}^{K-1}(u^{k}_{i\tau})^{2}$ (22)
[…]
III-B1 Collision Avoidance
The formulation of Problem 3 is agile and can be modified to achieve additional goals; for example, the cost function can be augmented with additional terms that ensure collision avoidance:
$J_{i}(y^{k}_{i},\bar{s}^{k}_{-i})=\sum_{\tau=1}^{K}(\ell^{k}_{i\tau}-1)^{2}+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2}\\ +G_{i}(s^{k}_{i},\bar{s}^{k}_{-i})+F_{i}^{\phi,\psi}(s^{k}_{i},\bar{s}^{k}_{-i}).$ (23)